Find the volume of the solid bounded by the elliptic paraboloids 2= 12 + 5y and 2 = 24 – 5.2 - 4
Answer #1

z utsy 2 = 24-55g Intersection of parabolicy xt5y = 24-55 yv but by =24 ta e cosa = esino - J J 24-50-5) - Entso) da. V=1/29

If you have any questions please let me know

Please give me up vote